help pls and explain it as well, thanks 😊
The Jones Family have a rectangular swimming pool. The pool is 5 metres longer than it is wide. There is a 2 metre wide strip of paving around the pool. The area of the paving is 54 square metres. Find the width of the swimming pool and give an equation that you use to solve the problem.

Found 2 solutions by josgarithmetic, mananth:
Answer by josgarithmetic(39800)
Answer by mananth(16949)

.
outer area - pool area = 54

(x+4)(x+9) - x(x+5)=54
x^2+13x+36-(x^2+5x)=54
x^2+13x+36-x^2-5x=54
8x+36=54
8x=54-36
8x=18
x=2.25
width = 2.5 m
length = 2.25+5 = 7.25m